OpenStudy (anonymous):

Compare and contrast simplifying the numeric expression 3 – (2 + 5) to the algebraic expression 3x – (2x + 5).

OpenStudy (anonymous):

if x=1 both are same

OpenStudy (phi):

They want you to see that in some ways, you treat them the same way. 3-(2+5)= 3 -2 -5 and 3x-(2x+5)= 3x -2x -5 You can then simply some more.